OpenStudy (anonymous):

Solve on the interval : [0,2pi] 4 csc x +6=-2 A. 7pi/6 , 11pi/6 B. pi/6 , 5pi/6 C. pi/3 , 5pi/3 D. 2pi/3 , 4pi/3

OpenStudy (anonymous):

Solve for x and determine the suitable x values for the expression and given interval. 4csc(x) = -4 csc(x) = -1 x = -π/2 x = 3π/2

OpenStudy (anonymous):

should be cscx = -2.... so sinx = -1/2 x = 7pi/6, 11pi/6